# ProtChatGPT: Towards Understanding Proteins with Large Language Models

### v1

@@ Abstract
Protein research is crucial in various fundamental disciplines, but understanding their intricate structure-function relationships remains challenging. Recent Large Language Models (LLMs) have made significant strides in comprehending task-specific knowledge, suggesting the potential for ChatGPT-like systems specialized in protein to facilitate basic research. In this work, we introduce ProtChatGPT, which aims at learning and understanding protein structures via natural languages. ProtChatGPT enables users to upload proteins, ask questions, and engage in interactive conversations to produce comprehensive answers. The system comprises protein encoders, a Protein-Language Pertaining Transformer (PLP-former), a projection adapter, and an LLM. The protein first undergoes protein encoders and PLP-former to produce protein embeddings, which are then projected by the adapter to conform with the LLM. The LLM finally combines user questions with projected embeddings to generate informative answers. Experiments show that ProtChatGPT can produce promising responses to proteins and their corresponding questions. We hope that ProtChatGPT could form the basis for further exploration and application in protein research. Code and our pre-trained model will be publicly available.
Protein Understanding, Large Language Models, ChatGPT-like system
1 Introduction
@@
Proteins, as essential molecular entities for life, hold paramount significance in biological processes. The comprehensive understanding of protein structure and function is of utmost importance for advancing research in the realms of biology and biomedicine. However, traditional protein research normally involves labor-intensive laboratory experiments and extensive literature reviews, which could be time-consuming and require specialized expertise in protein.
@@
Recently, Large Language Models (LLMs), e.g., ChatGPT (Radford et al., 2019), have prevailed in Natural Language Processing (NLP) (Devlin et al., 2018; Raffel et al., 2020; Touvron et al., 2023; Chowdhery et al., 2022). With superior language understanding and logical reasoning capabilities, these models can perform various intricate linguistic tasks such as question and answering (Q&A). Considering that protein sequences can be regarded as their own “natural language”, this intuitively motivates us to ride on LLMs’ coattails and customize them into protein research based on large-scale biological corpora, e.g., RCSB-PDB (Guo et al., 2023).
@@
Empirically, with the capabilities of LLMs specialized in protein, researchers can potentially achieve (1) Protein Understanding and Analysis by simplifying the retrieval of crucial information (e.g., structures, functions, interactions, mutations, and disease associations) about specific proteins for research; (2) Customized Protein Design by characterizing the patient’s unique protein structures to discover targeted drugs and further verify expected functions for healthcare.
@@
In this paper, we propose an AI-based protein chat system, named ProtChatGPT, to implement ChatGPT-like functionalities for the protein research field. ProtChatGPT works in a similar principle with natural language conversation systems. Users are allowed to upload protein 1D sequences or 3D structures (e.g., fasta or pdb files) and pose diverse related questions. Then, ProtChatGPT produces comprehensive responses in an interactive manner based on the questions. In this way, researchers can intuitively acquire valuable insights and interactively delve into the complexities of diverse proteins. Specifically, ProtChatGPT consists of four components: protein encoders, a Protein-Language Pre-training Transformer (PLP-former), a projection adapter and an LLM, as shown in Figure 1. First, we employ two pre-trained protein encoders to embed the 1D (i.e., by ESM-1b (Rives et al., 2021) ) and 3D structures (i.e., by ESM-IF1 (Hsu et al., 2022)), respectively. Then, to align the protein and language modalities, we propose the PLP-former. PLP-former extracts features from the output of the protein encoder, and learns the protein representations that are most relevant to the text description. Third, we use an adapter as an information transmitter to convert protein embeddings into protein prompts that can be interpreted by the LLM. Finally, the LLM combines user questions (i.e., question prompts) with the transmitted protein prompts to produce corresponding answers. We conduct experiments on protein understanding and design. Experimental results show the effectiveness of the proposed method. In summary, our contributions are as follows:
@@
  • We introduce PLP-former, a transformer-based module that aligns the protein with its corresponding description.
  • We propose a two-stage strategy that bootstraps protein-language pre-training from off-the-shelf pre-trained protein encoders and frozen large language models.
@@
Figure 1: An overview of the ProtChatGPT framework. The training process consists of two stages: (1) protein-description representation learning stage, and (2) protein-to-text generative learning stage. In the first stage, we enforce the PLP Transformer, a lightweight transformer with learnable query tokens, to extract features from a frozen protein 1D encoder, and learn the protein representation most relevant to the text description. The PLP-former takes sequence-description pairs as input, and outputs the learned tokens as the sequence embedding. In the second stage, we perform protein-to-text generative learning by connecting the output of the PLP Transformer as well as a frozen 3D encoder to a frozen LLM, and trains a multi-level adapter as an information bottleneck between two stages, such that its output protein representation can be interpreted by the LLMs. Finally, the LLM can produce descriptive answers given the question prompt and the multi-level (sequence and structure) protein prompt from the adapter.
2 Related Work
@@
Protein Representation Learning. Proteins are workhorses of the cell, which contain four distinct levels of structures carrying out their fundamental functions. Previous protein representation works seek to learn protein representations based on different levels of proteins. Considering protein sequences as language in life, several works (Madani et al., 2023; Notin et al., 2022) encode amino acid tokens using the Transformer model (Vaswani et al., 2017) to extract pairwise relationships among amino acids, and autoregressively recover protein sequences on extensive protein sequence databases. Alternatively, other sequence modeling methods(Lin et al., 2023; Meier et al., 2021; Rives et al., 2021; Rao et al., 2020; Vig et al., 2020) resort to use Masked Language Modeling (MLM) to develop attention patterns that correspond to the residue-residue contact map of the protein. Compared with sequence-based methods, structure-based methods (Gligorijević et al., 2021; OpenAI, 2023; Zhang et al., 2023) directly dictate protein functions and encode geometric information of proteins for topology-sensitive tasks such as molecule binding (Jin et al., 2021; Kong et al., 2022), protein interface analysis (Mahbub & Bayzid, 2022; Réau et al., 2023), and protein properties prediction (Zhang et al., 2022). In this paper, we aim to leverage these pre-trained Large Protein Models (LPMs) for high-quality embeddings without fine-tuning their network parameters.
Large Language Models (LLMs). Recently, Natural Language Processing (NLP) has witnessed significant advances due to the development of Large Language Models (LLMs) (Radford et al., 2019; Devlin et al., 2018) trained on an extensive, diverse corpus. Consequently, many multi-modal variants (Taylor et al., 2022; Alayrac et al., 2022; Jing et al., 2020; Wang et al., 2023; van Sonsbeek et al., 2023) based on LLMs have gained significant attention for the understanding of information in other modalities beyond text. For example, BLIP-2(Li et al., 2023) designs a Q-Former to align the visual features from the frozen visual encoder with large language models. FROMAGe (Koh et al., 2023) freezes the LLM and visual encoders, and fine-tunes linear mapping layers to achieve cross-modality interactions. Similarly, MedVQA (van Sonsbeek et al., 2023) employs a multi-layer perceptron (MLP) network that maps the extracted visual features from a frozen vision encoder to a set of learnable tokens, which develops an open-ended VQA for diagnoses and treatment decisions. Galactica (Taylor et al., 2022) explicitly models the protein sequences and SMILES with scientific literature, and enables the model to explain the properties of the sequences. ProteinChat (Guo et al., 2023) further models the protein structure with its corresponding descriptions in Protein Data Bank for the protein Q&A task. However, this method only considers the impact of protein structure on its function, while neglecting the complementarity between different levels of structure, such as amino acid sequences (primary structure) and geometric coordinates (tertiary structure). In this paper, we aim to adapt pre-trained general LLMs (e.g., Vicuna (Chiang et al., 2023)) for protein-specific ChatGPT-like tasks, which aligns multi-level protein features from LPMs with LLMs.
Vision-Language Pretraining (VLP). Data collected from different modalities generally offer distinct perspectives, frequently synergizing to yield a comprehensive understanding, enhancing the overall comprehension of the data. Vision-language pre-training (VLP) aims to learn multimodal foundation models, showing improved performance on various vision-and-language tasks (Radford et al., 2021a), Existing VLP methods can be roughly divided into representation learning-based and generative learning-based. Representation learning-based methods (Radford et al., 2021b; Jia et al., 2021; Yao et al., 2021; Li et al., 2022b, 2021) usually consider the image-text pairs as multi-modal views of the same semantics, and perform contrastive or multi-view learning for the alignment between multiple modalities. Generative learning-based methods  (Li et al., 2019; Lu et al., 2019; Chen et al., 2020; Li et al., 2020; Zhang et al., 2021; Wang et al., 2022; Zeng et al., 2021; Bao et al., 2022) aim to reconstruct the corrupted text (image) with the assistance of visual (text) modality through MLM-like objectives. For example, SimVLM (Wang et al., 2021) introduces a single prefix language modeling (PrefixLM) objective for exploiting large-scale weak supervision in VLP. CoCa (Yu et al., 2022) further verifies the representation ability of autoregressive language modeling in the vision-language domain. In this paper, we consider protein as a specialized biological language that encodes and communicates biological information through its amino acid sequences and interactions. Inspired by existing VLP methods (Li et al., 2023; Zhu et al., 2023), we first propose a representation learning-based Protein-Language Pre-training (PLP) framework to understand protein sequences via natural language instructions, and then design a generative learning-based multi-level adapter to generate the question-related answers combined with extra structural embeddings.
3 Methods
While demonstrating excellent performance in natural language tasks, LLM cannot directly facilitate protein question-answering tasks due to the modality gap between protein structures and biomedical texts. In order to bridge this gap, we introduce a protein-language pre-training strategy with two stages: (1) protein-description representation learning stage and (2) protein-to-text generative learning stage.
3.1 Architecture
As shown in Figure 1, ProtChatGPT consists of two pre-trained protein encoders, a trainable PLP-former, a trainable multi-level projection adapter and a pre-trained LLM decoder, working synergistically to provide protein-related insights.
Multi-Level Protein Encoders
First, we use a pre-trained sequence encoder to extract protein sequence features. Specifically, given a protein sequence with $N$ amino acids, the encoder produces the corresponding sequence embedding $\boldsymbol{E}_{seq}\in\mathbb{R}^{N\times C_{seq}}$, where $C_{seq}$ is the number of embedding channels. In our implementation, we use ESM-1b (Rives et al., 2021) as the 1D sequence encoder, where $C_{seq}$ = 768. Although ESM-1b is able to implicitly capture structural contact information, incorporating detailed 3D structures explicitly can be an effective way to model spatial interactions between residues. Therefore, we further propose to enhance the ESM-1b with a supplementary protein structure encoder ESM-IF1 (Hsu et al., 2022). Specifically, we select the feature from an intermediate layer as a complementary structure embedding $\boldsymbol{E}_{str}\in\mathbb{R}^{N\times C_{str}}$ where the number of embedding channels $C_{str}=512$, indicating the geometric protein knowledge. By acquiring such multi-level property information, the quality of protein embeddings can be further improved, considering that the protein properties studied in ESM-1b and ESM-IF1 can correlate with each other. Note that, both the sequence and structure encoder are frozen for efficient training.
@@ 3.1 Architecture
Figure 2: Illustrations of the PLP-Former and protein-language representation learning. PLP-Former consists of two transformer submodules with shared self-attention: (1) a text transformer that performs encoding and decoding of protein descriptions, and (2) a protein transformer that interacts with the frozen ESM-1b for sequence feature extraction. PLP-former is trained by jointly optimizing three pre-training objectives (dashed boxes) on sequence-description pairs.
Protein-Language Pretraining Transformer (PLP-former).
@@
As mentioned before, the key challenge of transferring LLMs to protein research lies in the modality gap between protein structures and biomedical texts. Despite the strong language generation and zero-shot transfer abilities of LLMs, directly retraining them end-to-end for protein specialization appears to be impractical due to the massive number of parameters and data requirements. Another alternative is fine-tuning the pre-trained parameters, but this often leads to catastrophic forgetting. Considering this trade-off, we propose a Protein-Language Pre-training Transformer (PLP-former) for efficient cross-modal alignment between protein and text, while remaining LLMs frozen during the training. Following existing vision-language works (Li et al., 2023; Zhu et al., 2023; Dai et al., 2023), we use the PLP-former to extract protein-related features from a frozen protein sequence encoder.
@@
As shown in Figure 2, the input of PLP-former consists of three parts: sequence embedding from ESM-1b, the corresponding description, and a set of learnable tokens. The learnable tokens $\boldsymbol{T}$ first perform mutual interactions via self-attention layers in the protein transformer. Specifically, the tokens $\boldsymbol{T}$ first acquire queries $\mathbf{Q}$, keys $\mathbf{K}$ and values $\mathbf{V}$ through three linear transformation matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$. It can be formulated as:
$\mathbf{Q}=\boldsymbol{E}_{str}^{l-1}\mathbf{W}_{q},\mathbf{K}=\boldsymbol{E}_{seq}^{l-1}\mathbf{W}_{k},\mathbf{V}=\boldsymbol{E}_{seq}^{l-1}\mathbf{W}_{v}.$ (1)
Next, the attention map $w$ is computed by taking the dot product of Q and K as:

### v2

# ProtChatGPT: Towards Understanding Proteins with Hybrid Representation and Large Language Models

@@ Abstract.
Protein research is crucial in various scientific disciplines, but understanding their intricate structure-function relationships remains challenging. Recent advancements in Large Language Models (LLMs) have significantly improved the comprehension of task-specific knowledge, suggesting the potential for specialized ChatGPT-like systems in protein research to aid fundamental investigations. In this work, we introduce ProtChatGPT, which aims to learn and understand protein structures using natural language. ProtChatGPT enables users to upload proteins, ask questions, and engage in interactive conversations to produce comprehensive answers. The system comprises multi-level protein encoding, protein-language alignment, and instruction tuning of LLMs. A protein first undergoes multiple protein encoders and PLP-former to produce multi-level hybrid protein embeddings, which are then aligned through a Protein Context Gating (PCG) module with contrastive learning, and projected by an adapter to conform with the LLM. The LLM finally combines user questions with projected protein embeddings to generate informative answers. Experiments show that ProtChatGPT can produce promising responses to proteins and the corresponding user questions. We hope that ProtChatGPT could form the basis for further exploration and application in protein research. Code and our pre-trained model will be publicly available.
Information retrieval, instruction tuning, large language models, protein understanding, question answering
ccs: Information systems Personalization
1. Introduction
@@
Proteins, as essential molecular entities for life, hold paramount significance in biological processes. The comprehensive understanding of protein structure and function is important for advancing research in the realms of biology and biomedicine. However, traditional protein research normally involves labor-intensive laboratory experiments and extensive literature reviews, which could be time-consuming and require specialized expertise in protein.
@@
Recently, Large Language Models (LLMs), e.g., ChatGPT (Radford et al., 2019), have prevailed in Natural Language Processing (NLP) (Raffel et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023). With superior language understanding and logical reasoning capabilities, these models can perform various intricate linguistic tasks such as question and answering (Q&A) (Zhu et al., 2023; Li et al., 2023). Considering that primary sequences can be regarded as protein’s own “natural language”, this intuitively motivates us to ride on LLMs’ coattails and customize them into protein research based on large-scale biological corpora, e.g., RCSB-PDB (Guo et al., 2023).
@@
Empirically, with the capabilities of LLMs specialized in protein, researchers can potentially achieve (1) Protein Understanding and Analysis by simplifying the retrieval of crucial information (e.g., structures, functions, interactions, mutations, and disease associations) about specific proteins for research; (2) Customized Protein Design by characterizing the unique protein structures of patients to discover targeted drugs and further verify expected functions for healthcare.
Figure 1. Overview of the ProtChatGPT framework. Our pipeline consists of three stages: (1) multi-level protein encoding, (2) multi-level protein-language alignment, and (3) instruction tuning with an external web corpus and protein features. First, we utilize three pre-trained frozen large protein encoders to acquire high-quality multi-level embeddings. In the second stage, we first enforce the PLP Transformer, a lightweight transformer with learnable query tokens, to learn the protein representation most relevant to the text description. The PLP-former takes the sequence embedding ${\bm{E}}_{seq}$, tokens, and protein descriptions as inputs and outputs the learned tokens as the selected embedding. The selected embedding is first aligned with the secondary structure embedding ${\bm{E}}_{sec}$ through protein context gating for aligned embedding ${\bm{E}}_{align}$, and then we adopt contrastive learning to enforce the tertiary structure embedding ${\bm{E}}_{ter}$ to further align with the joint representation ${\bm{E}}_{align}$. In the third stage, we perform protein-to-text generative learning by connecting the aligned ${\bm{E}}_{align}$ and ${\bm{E}}_{str}$ to an LLM decoder. Combined with the instruction pairs extracted from open protein-related literature, an adapter is further trained as the information bottleneck between protein embeddings and the LLM, such that its output can be interpreted by the language model. Finally, the LLM can produce descriptive answers given the question prompt and the multi-level protein prompt from the adapter.
@@
In this paper, we propose an AI-based protein chat system, named ProtChatGPT, to implement ChatGPT-like functionalities for the protein research field. ProtChatGPT works in a similar way to natural language conversation systems. Users are allowed to upload protein 1D sequences or 3D structures (e.g., fasta or pdb files) and pose diverse related questions. Then, ProtChatGPT produces comprehensive responses in an interactive manner based on the questions. In this way, researchers can intuitively acquire valuable insights and interactively delve into the complexities of diverse proteins. Specifically, ProtChatGPT consists of three stages: multi-level protein encoding, protein-language alignment, and instruction tuning of LLMs, as shown in Figure 1. First, we employ three pre-trained protein encoders to embed the primary sequence (i.e., by ESM-1b (Rives et al., 2021)), secondary structure (i.e., by NetSurfP (Klausen et al., 2019)), and tertiary structure (i.e., by ESM-IF1 (Hsu et al., 2022)), respectively. Next, to align the protein sequence and language modalities, we propose the PLP-former. PLP-former extracts features from the output of the protein encoder, and learns the protein representations that are most relevant to the text description. We also deploy a Protein Context Gating (PCG) module trained with contrastive loss further to align the sequence with other levels of protein embeddings. Third, we use an adapter as an information transmitter to convert protein embeddings into protein prompts that can be interpreted by the LLM. Finally, the LLM combines user questions (i.e., question prompts) with the transmitted protein prompts to produce corresponding answers. We conduct experiments on protein understanding and design. Experimental results show the effectiveness of the proposed method. In summary, our contributions are as follows:
@@
  • We introduce a progressive protein-language learning strategy for dynamically aligning protein from multi-level structures with natural language, including a transformer-based PLP-former and a Protein Context Gating (PCG) module based on contrastive learning.
  • As a minor contribution, a protein-specific instruction tuning dataset is built to facilitate the LLMs to better correspond to the user questions.
@@
2. Related Work
@@
Protein Representation Learning. Proteins are workhorses of the cell, which contain four distinct levels of structures carrying their fundamental functions. Previous protein representation works seek to learn protein representations based on different levels of structures. Considering protein sequences as language, several works (Madani et al., 2023; Notin et al., 2022) encode amino acid tokens using transformer model (Vaswani et al., 2017) to extract pairwise relationships among amino acids, and autoregressively recover protein sequences based on extensive protein sequence databases. Alternatively, other sequence modeling methods(Lin et al., 2023; Meier et al., 2021; Rives et al., 2021; Rao et al., 2020; Vig et al., 2020) resort to using Masked Language Modeling (MLM) to develop attention patterns that correspond to the residue-residue contact map of the protein. Compared with sequence-based methods, structure-based methods (Gligorijević et al., 2021; OpenAI, 2023; Zhang et al., 2023) directly encode geometric information of proteins for topology-sensitive tasks such as molecule binding (Jin et al., 2021; Kong et al., 2022), protein interface analysis (Mahbub and Bayzid, 2022; Réau et al., 2023), and protein properties prediction (Zhang et al., 2022a). In this paper, we aim to directly leverage these pre-trained Large Protein Models (LPMs) to acquire high-quality embeddings without fine-tuning their network parameters.
@@
Figure 2. Illustration of the PLP-former and protein-language representation learning. PLP-former consists of two transformer submodules with shared self-attention: (1) a text transformer that performs encoding and decoding of protein descriptions, and (2) a protein transformer that interacts with the frozen ESM-1b for sequence feature extraction. PLP-former is trained by jointly optimizing three pretraining objectives (dashed boxes) on sequence-description pairs.
Large Language Models (LLMs). Recently, Natural Language Processing (NLP) has witnessed significant advances due to the development of LLMs (Radford et al., 2019; Devlin et al., 2018) trained on an extensive, diverse dataset. Consequently, many multi-modal variants (Taylor et al., 2022; Alayrac et al., 2022; Jing et al., 2020; Wang et al., 2023; van Sonsbeek et al., 2023) based on LLMs have gained significant attention for the understanding of information in other modalities beyond text. For example, BLIP-2(Li et al., 2023) designs a Q-Former to align the visual features from the frozen visual encoder with large language models. FROMAGe (Koh et al., 2023) freezes the LLM and visual encoders, and fine-tunes linear mapping layers to achieve cross-modality interactions. For protein understanding, MedVQA (van Sonsbeek et al., 2023) employs a Multi-Layer Perceptron (MLP) network that maps the extracted visual features from a frozen vision encoder to a set of learnable tokens, which develops an open-ended VQA for diagnoses and treatment decisions. Galactica (Taylor et al., 2022) explicitly models the protein sequences and SMILES with scientific literature, and enables the model to explain the properties of the sequences. ProteinChat (Guo et al., 2023) further models the protein structure with its corresponding descriptions in Protein Data Bank (PDB) for the protein Q&A task. However, this method only considers the impact of protein structure on its function, while neglecting the complementarity between different levels of structure, such as amino acid sequences (primary structure) and geometric coordinates (tertiary structure). In this paper, we aim to adapt pre-trained general LLMs (e.g., Vicuna (Chiang et al., 2023) and LL A MA3) for protein-specific ChatGPT-like tasks, which aligns multi-level protein features from LPMs with LLMs.
3. Proposed Method
While demonstrating excellent performance in natural language tasks, LLM cannot directly facilitate protein question-answering tasks due to the modality gap between protein structures and biomedical descriptions. As shown in Figure 1, in order to bridge this gap, we introduce a progressive protein-language learning framework with three stages: (1) multi-level protein encoding, (2) multi-level protein-language alignment, and (3) instruction tuning with protein features.
3.1. Multi-Level Protein Encoding
In this paper, we leverage three pre-trained Large Protein Models (LPMs) exclusively trained on different levels of protein structures, working synergistically to provide high-quality hybrid embeddings without fine-tuning their network parameters.
First, we use a pre-trained sequence encoder to extract protein sequence features. Specifically, given a protein sequence with ${n}$ amino acids, the encoder produces the corresponding sequence embedding ${\bm{E}}_{seq}\in\mathbb{R}^{{n}\times{c}_{seq}}$, where ${c}_{seq}$ is the number of embedding channels. In our implementation, we use ESM-1b (Rives et al., 2021) as the sequence encoder, where ${c}_{seq}$ = 768. Although ESM-1b is able to implicitly capture structural contact information, incorporating detailed structure (secondary and tertiary) information explicitly can be an effective way to model spatial interactions between residues. Therefore, we propose to enhance the ESM-1b with a supplementary protein secondary structure encoder NetSurfP (Klausen et al., 2019). Specifically, this encoder consists of two parallel convolution layers and an identity layer, whose outputs are concatenated in the feature dimension and fed to a two-layer bidirectional LSTM containing 1024 hidden units. The output is then projected to an 8-dimensional feature vector at each position as the secondary structure embedding ${\bm{E}}_{sec}\in\mathbb{R}^{{n}\times 8}$. Moreover, we further introduce the ESM-IF1 (Hsu et al., 2022) encoder for protein tertiary structure representation. Specifically, we select the feature from an intermediate layer as the tertiary structure embedding ${\bm{E}}_{ter}\in\mathbb{R}^{{n}\times{c}_{ter}}$ where the number of embedding channels ${c}_{ter}=512$, indicating the geometric protein knowledge.
By acquiring such multi-level hybrid property information, the quality of protein embeddings can be further improved, considering that the protein properties studied in ESM-1b and ESM-IF1 can correlate with each other. Note that all the employed encoders are frozen for efficient training.
3.2. Multi-Level Protein-Language Alignment
3.2.1. Protein-Language Pretraining (PLP)
@@
As mentioned before, the key challenge of transferring LLMs to protein research lies in the modality gap between protein structures and biomedical texts. Despite the strong language generation and zero-shot transfer abilities of LLMs, directly retraining them end-to-end for protein specialization appears to be impractical due to the massive number of parameters and data requirements. Another alternative is fine-tuning the pre-trained parameters, but this often leads to catastrophic forgetting. Considering this trade-off, we propose a Protein-Language Pretraining Transformer (PLP-former) for efficient cross-modal alignment between protein and text, while remaining LLMs frozen during the training. Following existing Vision-Language Pretraining (VLP) works (Li et al., 2023; Zhu et al., 2023; Dai et al., 2023), we introduce the PLP-former to extract protein-related features from a frozen protein sequence encoder. More related works on VLP are also given in Appendix D.
@@
As shown in Figure 2, PLP-former consists of three inputs: initial sequence embedding ${\bm{E}}_{seq}\in\mathbb{R}^{{n}\times 768}$ from ESM-1b, a set of learnable tokens ${\bm{T}}\in\mathbb{R}^{32\times 768}$, and the corresponding description. The learnable tokens ${\bm{T}}$ first perform mutual interactions via self-attention layers in the protein transformer. Specifically, the tokens ${\bm{T}}$ first acquire queries ${\bm{Q}}$, keys ${\bm{K}}$ and values ${\bm{V}}$ through three linear transformation matrices ${\bm{W}}_{q}$, ${\bm{W}}_{k}$ and ${\bm{W}}_{v}$. It can be formulated as:
(1) ${\bm{Q}}={\bm{T}}{\bm{W}}_{q},{\bm{K}}={\bm{T}}{\bm{W}}_{k},{\bm{V}}={\bm{T}}{\bm{W}}_{v}.$
Next, the attention map ${\bm{M}}$ is computed through $\mathrm{softmax}(\frac{{\bm{Q}}{\bm{K}}^{T}}{\sqrt{{d}_{k}}})$ where ${d}_{k}$ represents the dimensionality of the keys, softmax is the softmax activation function. The refined tokens ${\bm{T}}^{\prime}\in\mathbb{R}^{32\times 768}$ after self-attention can be calculated as ${\bm{T}}^{\prime}={\bm{W}}_{o}{\bm{M}}{\bm{V}}$ where ${\bm{W}}_{o}$ is the projection matrix for output. Given the sequence features from ESM-1b as ${\bm{E}}_{seq}\in\mathbb{R}^{{n}\times 768}$, tokens ${\bm{T}}^{\prime}\in\mathbb{R}^{32\times 768}$ then interact with language embedding through cross-attention layers, which can be formulated as:
(2) ${\bm{T}}^{\prime\prime}={\bm{W}}_{o}^{\prime}\left(~{}\mathrm{softmax}\left(\frac{{\bm{T}}^{\prime}{\bm{W}}_{q}^{\prime}({\bm{E}}_{seq}{\bm{W}}_{k}^{\prime})^{T}}{\sqrt{d_{k}}}\right)\left({\bm{E}}_{seq}{\bm{W}}_{v}^{\prime}\right)\right),$
where ${\bm{T}}^{\prime\prime}\in\mathbb{R}^{32\times 768}$ represents the refined tokens after cross-attention, ${\bm{W}}_{q}^{\prime}$, ${\bm{W}}_{k}^{\prime}$, ${\bm{W}}_{v}^{\prime}$ and ${\bm{W}}_{o}^{\prime}$ are a new set of learning transformation matrices. Additionally, benefiting from the text transformer, tokens can further interact with the textual descriptions through the same self-attention layers. Finally, the PLP-former produces the output ${\bm{E}}_{seq}$ after a linear feed-forward layer (Vaswani et al., 2017). The size of selected ${\bm{E}}_{seq}$ ($32\times 768$) is much smaller than the size of frozen sequence features (${n}\times 768$) as the length of protein ${n}$ usually numbers over hundreds. Depending on the pretraining task, we implement distinct attention masking strategies (Li et al., 2023) within the self-attention block to regulate the token-text interaction. The training details of PLP-former are given in Appendix E. In this way, PLP-former can effectively select the most useful information for the LLM while removing irrelevant protein information. This reduces the burden of the LLM to learn protein-language alignment while mitigating the catastrophic forgetting problem, as the PLP-former establishes a fundamental correlation between the primary structure and its descriptive text before introducing more complex structural data.
3.2.2. Multi-Level Protein Alignment
After multi-level encoding, the next step involves aligning proteins with these multi-level features to harness the full potential of their hybrid structural representations. In this work, we introduce a progressive alignment strategy across different levels.
Protein Context Gating (PCG)
Initially, we utilize a Protein Context Gating (PCG) module for the feature aggregation between the sequence embedding ${\bm{E}}_{seq}\in\mathbb{R}^{32\times 768}$ and secondary structure embedding ${\bm{E}}_{sec}\in\mathbb{R}^{{n}\times 8}$. Our goal is to transform ${\bm{E}}_{sec}$ into dynamic gating maps, which effectively allows the PCG module to highlight or suppress ${\bm{E}}_{seq}$ based on the relevant secondary structural information. The formulation of PCG can be written as:
(3) ${\bm{E}}_{align}=\sigma({\bm{W}}_{sec}{\bm{E}}_{sec}+{\bm{b}})\odot{\bm{E}}_{seq},$
where $\odot$ represents the element-wise multiplication, $\sigma$ denotes the sigmoid activation function that normalizes the gating values to the range [0,1]. A convolution layer is first deployed on ${\bm{E}}_{sec}\in\mathbb{R}^{n\times 8}$ to project the dimension to $32\times 8$. ${\bm{W}}_{sec}\in\mathbb{R}^{8\times 768}$ is a transformation matrix, ${\bm{b}}\in\mathbb{R}^{1\times 768}$ is the bias to acquire the aligned embedding ${\bm{E}}_{align}\in\mathbb{R}^{32\times 768}$.
Contrastive Learning
Subsequently, we employ a contrastive loss to achieve alignment with the tertiary structure ${\bm{E}}_{ter}$, which can be formulated as:
(4) $\mathcal{L}=-\log\frac{\exp({\bm{E}}_{align}\cdot{\bm{E}}_{ter_{+}}/\tau)}{\sum_{i=0}^{k}\exp({\bm{E}}_{align}\cdot{\bm{E}}_{ter_{i}}/\tau)},$
where ${\bm{E}}_{ter_{+}}$ is the positive sample corresponding to $E_{align}$, $\cdot$ represents the dot product, $\tau$ represents a temperature hyper-parameter modulating the learning sensitivity. We also added a convolutional layer on ${\bm{E}}_{ter}\in\mathbb{R}^{n\times 512}$ to match the dimensions with ${\bm{E}}_{align}\in\mathbb{R}^{32\times 768}$. By introducing the tertiary structural information, this contrastive mechanism allows for dynamic protein feature alignment and optimizes the representation for tasks requiring high discernment between similar protein sequences, such as homologous protein analysis.
3.3. Instruction Tuning with Protein Features
3.3.1. Projection Adapter
During the final instruction tuning stage, we further design a multi-level projection adapter to harvest the generative language capability of LLMs for protein-to-text generative learning. The adapter takes the pre-aligned embeddings ${\bm{E}}_{align}$ and ${\bm{E}}_{ter}$ from ESM-IF1 as inputs, and acts as an information bottleneck to the LLM decoder, such that its output protein representation can be interpreted by the LLM. In practice, we use two individual Fully-Connected (FC) layers to linearly project the output protein embeddings into the same dimension as the question embedding of the LLM. They function as soft protein prompts that condition the LLM on protein representation from different levels.
3.3.2. Protein-Text Generation
Finally, the projected protein prompts are prepended to the question prompts (text embeddings of user questions) through concatenation. In implementation, we separately deploy two LLMs Vicuna-13B (Chiang et al., 2023) and LL A MA3. We employ the LLM decoder to model the conditional generation probability $p_{\theta}({\bm{t}}_{i}|{\bm{t}}_{<i})$ in the language model, where ${\bm{t}}_{i}$ is the generated tokens, ${\bm{t}}_{<i}$ represents the context vector of input tokens (i.e., protein embeddings along with user questions). To enhance the model training with protein-text pairs, we utilize a specialized token prompt as:
$\mathbf{Q}{:}&<{\text{Protein}}><\text{ProteinPrompts}></\text{Protein}>\\ &<\text{QuestionPrompts}>\\ \mathbf{A}{:}&<\text{Descriptions}>\end{aligned}$